OpenStudy (anonymous):

use the continuity of the exponential function to show that lim x->0 (1 + tx)^(1/x) = e^t @Mathematics

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (1+tx)^{1/x}=e ^{t}\]

OpenStudy (anonymous):

Take the natural log of both sides, direct substitution, then solve for t.

OpenStudy (anonymous):

what happens when I take the log of a limit?

OpenStudy (anonymous):

why am I trying to solve for t?

OpenStudy (anonymous):

Bahh it keeps bugging out!!!!

OpenStudy (anonymous):

I don't even know how to begin going about solving this

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (1+tx)^{\frac{1}{x}}\] Now let k=1/x \[\lim_{k \rightarrow \infty } \left ( 1+\frac{t}{k}\right )^{k}\] Now let k=ta \[\lim_{a \rightarrow \infty } \left ( 1+\frac{1}{a}\right )^{at} = e^{t}\]